OpenStudy (anonymous):

I GIVE BEST MEDAL IN THE WORLD A gym offers regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships? 125 140 235 295

OpenStudy (anonymous):

295

OpenStudy (anonymous):

can you explain to me why its 295

OpenStudy (anonymous):

yes, regular memberships is x regular is y 80x+60y=31100 x+y=420 or you could just do 295*80=23600 420-295=125 125*60=7500 23600+7500=31100 so the answer is 295